# The Poisson Midpoint Method for Langevin Dynamics: Provably Efficient Discretization for Diffusion Models

### v2

@@ F.2 Bounding the Bias:
$\leq C\frac{\alpha^{5}L^{4}N_{t}^{2}}{K\gamma}\sup_{0\leq i\leq K-1}\|\hat{\tilde{U}}_{tK+i}-\tilde{U}_{tK}\|^{2}$ (56)
@@
$\leq C\alpha^{7}L^{3}\gamma T(d+\log K)+\frac{C\alpha^{6}L^{4}}{\gamma^{2}}\left[\mathbb{E}\|\tilde{V}_{0}\|^{2}+\mathbb{E}(F(\Psi_{0})-F(\Psi_{T}))^{+}+1\right]$ (57)
@@ F.3 Bounding the Variance:
$\leq\frac{CK\alpha L^{2}}{\gamma}\left[\alpha^{2}\|\tilde{V}_{tK}\|^{2}+\alpha^{4}\|\nabla F(\tilde{U}_{tK})\|^{2}+M_{t,K}^{2}\right]$ (58)
@@
$\quad+C_{p}\frac{K^{p+1}\alpha^{3p}L^{2p}T}{\gamma^{p}}\left(\frac{\gamma^{2p}}{L^{p}}+(\gamma\alpha T)^{p-1}\right)(d+\log K)^{p}$ (59)
@@ F.4 Finishing the Proof
$\quad+\frac{K^{7}\alpha^{27}L^{14}T^{7}}{\gamma}(d+\log K)^{7}$ (60)
@@ Appendix G Convergence Under Isoperimetry Assumptions
$\mathbb{E}_{\pi^{\star}}g^{2}\log g^{2}-\mathbb{E}_{\pi^{\star}}g^{2}\log\mathbb{E}_{\pi^{\star}}g^{2}\leq\frac{2}{\lambda}\mathbb{E}_{\pi^{\star}}\|\nabla g\|^{2}$
@@
$\mathbb{E}_{\pi^{\star}}g^{2}-(\mathbb{E}_{\pi^{\star}}g)^{2}\leq\frac{1}{\lambda}\mathbb{E}_{\pi^{\star}}\|\nabla g\|^{2}$
@@
$\mathsf{var}(\langle X,v\rangle)\leq\frac{1}{\lambda}$
@@
$\leq\frac{1}{2d}\mathbb{E}\|X-X^{\prime}\|^{2}=\frac{\mathsf{Tr}(\Sigma)}{d}\leq\frac{1}{\lambda}$ (61)
@@
$\mathsf{TV}(\mathsf{Law}(X^{P}_{T}),\pi^{\star})\leq\sqrt{2\mathsf{KL}\left(\mathsf{Law}(X^{P}_{T})\boldsymbol{\bigr{|}\bigr{|}}\mathsf{Law}(X_{TK})\right)}+\sqrt{2\mathsf{KL}\left(\mathsf{Law}(X_{TK})\boldsymbol{\bigr{|}\bigr{|}}\pi^{\star}\right)}\leq\epsilon$ (62)
@@
$\mathsf{TV}(\mathsf{Law}(U_{N}),\pi^{\star})\leq\epsilon$
@@
$\mathsf{TV}(\mathsf{Law}(X_{T}^{(P)}),\pi^{\star})\leq\epsilon$
@@ Appendix H Some Technical Results
$\sum_{j=1}^{T}\mathbb{E}\|a_{j}\|^{2p}\leq\frac{2^{p}}{N}\sum_{k=1}^{\frac{T}{N}}\sum_{j\in\mathcal{T}_{k}}\sum_{\begin{subarray}{c}j^{\prime}\in\mathcal{T}_{k}\\ j^{\prime}>j\end{subarray}}\mathbb{E}\|a_{j}-a_{j^{\prime}}\|^{2p}+\left(\frac{2}{N}\right)^{p-1}\mathbb{E}(\sum_{j=1}^{T}\|a_{j}\|^{2})^{p}$
@@
$\mathbb{E}|\sum_{s=1}^{T}\langle g_{s},Z_{s}\rangle|^{p}\leq C_{p}\sqrt{\mathbb{E}(\sum_{s=1}^{T}\|g_{s}\|^{2})^{p}}$
@@
[…]
@@ I.1 Proof of Lemma 12
    $\leq\alpha N_{t}L\sup_{0\leq j\leq K-1}\|\hat{\tilde{X}}_{tK+j}-\tilde{X}_{tK}\|$ (63)
@@ I.2 Proof of Lemma 13
Where $\bar{Y}_{t}=\frac{1}{\sqrt{K}}\sum_{i=0}^{K-1}Z_{tK+i}$ and $\Delta_{t}=\sum_{i=0}^{K-1}H_{i}(b(\hat{\tilde{X}}_{tK+i})-b(\tilde{X}_{tK}))$
@@
$=\alpha(s-t)b(\tilde{X}_{tK})+\sum_{h=t}^{s-1}\alpha(b(\tilde{X}_{hK})-b(\tilde{X}_{tK}))+\alpha\Delta_{h}+\sqrt{2\alpha}\tilde{Y}_{h}$ (64)
@@
$\quad+C_{p}\sqrt{2\alpha d(s-t)}$ (65)
@@
$\leq LC_{p}\alpha\|b(\tilde{X}_{tK})\|+L^{2}C_{p}\alpha\mathcal{M}_{p}(\tilde{X}_{hK}-\tilde{X}_{tK})+LC_{p}\sqrt{\alpha d}$ (66)
@@
$\quad+C_{p}(L\alpha^{\tfrac{3}{2}}\sqrt{d}(s-t)+\sqrt{2\alpha d(s-t)})$ (67)
@@
$\quad+C_{p}(L\sqrt{d}\alpha^{\tfrac{3}{2}}(s-t)+\sqrt{2\alpha d(s-t)})$ (68)
@@
$\sup_{t\leq h\leq s}\mathcal{M}_{p}(\tilde{X}_{hK}-\tilde{X}_{tK})$ $\leq 3\alpha L(s-t)\frac{\|b(\tilde{X}_{tK})\|}{L}+C_{p}\sqrt{d\alpha(s-t)}$
@@ I.3 Proof of Lemma 26
$=\frac{2}{N}\sum_{j\in\mathcal{T}_{k}}\sum_{j^{\prime}>j}\mathbb{E}\|a_{j}-a_{j^{\prime}}\|^{2p}$ (71)
@@ I.4 Proof of Lemma 14
$\leq C_{p}L^{2p}(\alpha d)^{p}N^{p+1}+C_{p}(\alpha LN)^{2p}\sum_{j\in\mathcal{T}_{k}}\mathbb{E}\|a_{j}\|^{2p}$ (72)
@@ I.5 Proof of Lemma 15
$F(\tilde{X}_{(t+1)K})-F(\tilde{X}_{tK})\leq\langle\nabla F(\tilde{X}_{tK}),\tilde{X}_{(t+1)K}-\tilde{X}_{tK}\rangle+\frac{L}{2}\|\tilde{X}_{(t+1)K}-\tilde{X}_{tK}\|^{2}$ (75)
@@
Where $\bar{Y}_{t}=\frac{1}{\sqrt{K}}\sum_{i=0}^{K-1}Z_{tK+i}$ and $\Delta_{t}=-\sum_{i=0}^{K-1}H_{t,i}(\nabla F(\hat{\tilde{X}}_{tK+i})-\nabla F(\tilde{X}_{tK}))$
Thus, summing Equation (75) from $t=0$ to $t=T-1$, we conclude:
@@
$\leq\sum_{t=0}^{T-1}-\tfrac{\alpha}{4}\|\nabla F(\tilde{X}_{tK})\|^{2}+3\alpha\|\Delta_{t}\|^{2}+3L\alpha\|\bar{Y}_{t}\|^{2}+\sqrt{2\alpha}\langle\nabla F(\tilde{X}_{tK}),\bar{Y}_{t}\rangle$ (76)
@@
$\quad+\frac{C_{p}}{\alpha^{\frac{p}{2}}}\bigr{|}\sum_{t=0}^{T-1}\langle\nabla F(\tilde{X}_{tK}),\bar{Y}_{t}\rangle\bigr{|}^{p}$ (77)
@@
$\leq\frac{\kappa C_{p}}{2\alpha^{p}}+\frac{C_{p}}{2\kappa}\mathbb{E}(\sum_{t=0}^{T-1}\|\nabla F(\tilde{X}_{tK})\|^{2})^{p}$ (78)
@@
$\leq C_{p}T^{p-1}\sum_{t=0}^{T-1}L^{2p}\mathbb{E}(\sup_{0\leq i\leq K-1}\|\hat{\tilde{X}}_{tK+i}-\tilde{X}_{tK}\|)^{2p}$ (79)
In the first step, we have used Jensen’s inequality. In the second step we have used the fact that $\nabla F$ is $L$-Lipshcitz and that $N_{t}$ is independent of $\hat{\tilde{X}}_{tK+i}$ for $10\leq i\leq K-$ 1. In the last step, we have used the fact that $\mathbb{E}(N_{t})^{2p}\leq C_{p}$ for some constant $C_{p}$ (Lemma 28).
@@
$\mathbb{E}\sup_{0\leq i\leq k}\|\hat{\tilde{X}}_{tK+i}-\tilde{X}_{tK}\|^{2p}\leq C_{p}\alpha^{2p}\mathbb{E}\|\nabla F(\tilde{X}_{tK})\|^{2p}+C_{p}\alpha^{p}d^{p}$
@@
$\quad+C_{p}L^{p}d^{p}T^{p}+\frac{C_{p}}{\alpha^{p}}$ (80)
@@ I.8 Proof of Lemma 21
    Using the scaling relations (Section 2.1), it is easy to show that for any $i$ such that $10\leq i\leq K-$ 1, $\sum_{j=0}^{i-1}\Pi A_{\tfrac{\alpha(i-1-j)}{K}}\Gamma_{\tfrac{\alpha}{K}}Z_{tK+j}$ is a Gaussian with covariance matrix $\Sigma_{i}$ such that $\Sigma_{i}\preceq\frac{2\exp(\gamma\alpha)\gamma\alpha^{3}}{3}\mathbf{I}_{d}$.
    Applying Gaussian concentration for $\sum_{j=0}^{i-1}\Pi A_{\tfrac{\alpha(i-1-j)}{K}}\Gamma_{\tfrac{\alpha}{K}}Z_{tK+j}$ along with the union bound over all $10\leq i\leq K-$ 1, we conclude:
@@ I.9 Proof of Lemma 22
    In the second step, we have use the fact that $|\frac{2g(1-g)}{\gamma}\langle\nabla F(\tilde{U}_{tK}),\tilde{V}_{tK}\rangle|\leq g(1-g)\frac{\|\nabla F(\tilde{U}_{tK})\|^{2}}{\gamma^{2}}+g(1-g)\|\tilde{V}_{tK}\|^{2}$. In the third step, we have used the fact that: $2g\langle\tilde{V}_{tK},(\mathbf{I}-\Pi)\alpha\Delta_{t}\rangle\leq(\sqrt{g}-g)\|\tilde{V}_{tK}\|^{2}+\frac{\|(\mathbf{I}-\Pi)\alpha\Delta_{t}\|^{2}}{1-\sqrt{g}}$, $-\frac{2(1-g)}{\gamma}\langle\nabla F(\tilde{U}_{tK}),(\mathbf{I}-\Pi)\alpha\Delta_{t}\rangle\leq\frac{1-g}{\gamma^{2}}\|\nabla F(\tilde{U}_{tK})\|^{2}+(1-g)\|(\mathbf{I}-\Pi)\alpha\Delta\|^{2}$ and $\|(\mathbf{I}-\Pi)(\alpha\Delta_{t}+\Gamma_{\alpha}\tilde{Y}_{t})\|^{2}\leq 2\|(\mathbf{I}-\Pi)(\alpha\Delta_{t})\|^{2}+2\|(\mathbf{I}-\Pi)\Gamma_{\alpha}\tilde{Y}_{t}\|^{2}$.
@@
    We now use the fact that for any two random variables $X,Y$ such that $\mathbb{E}\|X\|^{p},\mathbb{E}\|Y\|^{p}<\infty$, $[\mathbb{E}\|X+Y\|^{p}]^{\frac{1}{p}}\leq[\mathbb{E}\|X\|^{p}]^{\frac{1}{p}}+[\mathbb{E}\|Y\|^{p}]^{\frac{1}{p}}$. Using the bounds established above and applying them to Equation (91), we conclude:
@@
    Summing item 2 of Lemma 21 over $t=0$ to $t=T-1$ and applying the triangle inequality for $p$-th moments, we have:

### v3

@@ F.2 Bounding the Bias:
$B_{tK+i}:=\Gamma_{\tfrac{\alpha}{K}}^{-1}G_{\tfrac{\alpha}{K}}[b(\hat{\tilde{X}}_{tK+i})-b(\tilde{X}_{tK+i})].$
@@
$\leq C\frac{\alpha^{5}L^{4}N_{t}^{2}}{K\gamma}\sup_{0\leq i\leq K-1}\|\hat{\tilde{U}}_{tK+i}-\tilde{U}_{tK}\|^{2}.$ (56)
@@
$\leq C\alpha^{7}L^{3}\gamma T(d+\log K)+\frac{C\alpha^{6}L^{4}}{\gamma^{2}}\left[\mathbb{E}\|\tilde{V}_{0}\|^{2}+\mathbb{E}(F(\Psi_{0})-F(\Psi_{T}))^{+}+1\right].$ (57)
@@ F.3 Bounding the Variance:
$\leq\frac{CK\alpha L^{2}}{\gamma}\left[\alpha^{2}\|\tilde{V}_{tK}\|^{2}+\alpha^{4}\|\nabla F(\tilde{U}_{tK})\|^{2}+M_{t,K}^{2}\right].$ (58)
@@
$\quad+C_{p}\frac{K^{p+1}\alpha^{3p}L^{2p}T}{\gamma^{p}}\left(\frac{\gamma^{2p}}{L^{p}}+(\gamma\alpha T)^{p-1}\right)(d+\log K)^{p}.$ (59)
@@ F.4 Finishing the Proof
$\quad+\frac{K^{7}\alpha^{27}L^{14}T^{7}}{\gamma}(d+\log K)^{7}.$ (60)
@@ Appendix G Convergence Under Isoperimetry Assumptions
$\mathbb{E}_{\pi^{\star}}g^{2}\log g^{2}-\mathbb{E}_{\pi^{\star}}g^{2}\log\mathbb{E}_{\pi^{\star}}g^{2}\leq\frac{2}{\lambda}\mathbb{E}_{\pi^{\star}}\|\nabla g\|^{2}.$
@@
$\mathbb{E}_{\pi^{\star}}g^{2}-(\mathbb{E}_{\pi^{\star}}g)^{2}\leq\frac{1}{\lambda}\mathbb{E}_{\pi^{\star}}\|\nabla g\|^{2}.$
@@
$\mathsf{var}(\langle X,v\rangle)\leq\frac{1}{\lambda}.$
@@
$\leq\frac{1}{2d}\mathbb{E}\|X-X^{\prime}\|^{2}=\frac{\mathsf{Tr}(\Sigma)}{d}\leq\frac{1}{\lambda}.$ (61)
@@
$\mathsf{TV}(\mathsf{Law}(X^{P}_{T}),\pi^{\star})\leq\sqrt{2\mathsf{KL}\left(\mathsf{Law}(X^{P}_{T})\boldsymbol{\bigr|\bigr|}\mathsf{Law}(X_{TK})\right)}+\sqrt{2\mathsf{KL}\left(\mathsf{Law}(X_{TK})\boldsymbol{\bigr|\bigr|}\pi^{\star}\right)}\leq\epsilon.$ (62)
@@
$\mathsf{TV}(\mathsf{Law}(U_{N}),\pi^{\star})\leq\epsilon.$
@@
$\mathsf{TV}(\mathsf{Law}(X_{T}^{(P)}),\pi^{\star})\leq\epsilon.$
@@ Appendix H Some Technical Results
$\sum_{j=1}^{T}\mathbb{E}\|a_{j}\|^{2p}\leq\frac{2^{p}}{N}\sum_{k=1}^{\frac{T}{N}}\sum_{j\in\mathcal{T}_{k}}\sum_{\begin{subarray}{c}j^{\prime}\in\mathcal{T}_{k}\\ j^{\prime}>j\end{subarray}}\mathbb{E}\|a_{j}-a_{j^{\prime}}\|^{2p}+\left(\frac{2}{N}\right)^{p-1}\mathbb{E}(\sum_{j=1}^{T}\|a_{j}\|^{2})^{p}.$
@@
$\mathbb{E}|\sum_{s=1}^{T}\langle g_{s},Z_{s}\rangle|^{p}\leq C_{p}\sqrt{\mathbb{E}(\sum_{s=1}^{T}\|g_{s}\|^{2})^{p}}.$
@@
Now, for any variable $X$, $\mathbb{P}(X^{+}>h)=\mathbb{P}(X>h)$ whenever $h\geq 0$. Now using the fact that: $\mathbb{E}X\leq\mathbb{E}X^{+}=\int_{0}^{\infty}\mathbb{P}(X^{+}>h)dh$, and taking $X=|\sum_{s=1}^{T}\langle g_{s},Z_{s}\rangle|^{p}-C_{p}\lambda^{p}(\sum_{s=1}^{T}\|g_{s}\|^{2})^{p}$:
$\mathbb{E}|\sum_{s=1}^{T}\langle g_{s},Z_{s}\rangle|^{p}\leq C_{p}\lambda^{p}\mathbb{E}(\sum_{s=1}^{T}\|g_{s}\|^{2})^{p}+\int_{0}^{\infty}2\exp(-c_{p}\lambda h^{\tfrac{1}{p}})dh\leq C_{p}\lambda^{p}\mathbb{E}(\sum_{s=1}^{T}\|g_{s}\|^{2})^{p}+\frac{C_{p}}{\lambda^{p}}$
@@ I.1 Proof of Lemma 12
    $\leq\alpha N_{t}L\sup_{0\leq j\leq K-1}\|\hat{\tilde{X}}_{tK+j}-\tilde{X}_{tK}\|.$ (63)
@@ I.2 Proof of Lemma 13
Here $\bar{Y}_{t}=\frac{1}{\sqrt{K}}\sum_{i=0}^{K-1}Z_{tK+i}$ and $\Delta_{t}=\sum_{i=0}^{K-1}H_{i}(b(\hat{\tilde{X}}_{tK+i})-b(\tilde{X}_{tK}))$.
@@
$=\alpha(s-t)b(\tilde{X}_{tK})+\sum_{h=t}^{s-1}\alpha(b(\tilde{X}_{hK})-b(\tilde{X}_{tK}))+\alpha\Delta_{h}+\sqrt{2\alpha}\tilde{Y}_{h}.$ (64)
@@
$\quad+C_{p}\sqrt{2\alpha d(s-t)}.$ (65)
@@
$\leq LC_{p}\alpha\|b(\tilde{X}_{tK})\|+L^{2}C_{p}\alpha\mathcal{M}_{p}(\tilde{X}_{hK}-\tilde{X}_{tK})+LC_{p}\sqrt{\alpha d}.$ (66)
@@
$\quad+C_{p}(L\alpha^{\tfrac{3}{2}}\sqrt{d}(s-t)+\sqrt{2\alpha d(s-t)}).$ (67)
@@
$\quad+C_{p}(L\sqrt{d}\alpha^{\tfrac{3}{2}}(s-t)+\sqrt{2\alpha d(s-t)}).$ (68)
@@
$\sup_{t\leq h\leq s}\mathcal{M}_{p}(\tilde{X}_{hK}-\tilde{X}_{tK})$ $\leq 3\alpha L(s-t)\frac{\|b(\tilde{X}_{tK})\|}{L}+C_{p}\sqrt{d\alpha(s-t)}.$
@@ I.3 Proof of Lemma 26
$=\frac{2}{N}\sum_{j\in\mathcal{T}_{k}}\sum_{j^{\prime}>j}\mathbb{E}\|a_{j}-a_{j^{\prime}}\|^{2p}.$ (71)
@@ I.4 Proof of Lemma 14
$\leq C_{p}L^{2p}(\alpha d)^{p}N^{p+1}+C_{p}(\alpha LN)^{2p}\sum_{j\in\mathcal{T}_{k}}\mathbb{E}\|a_{j}\|^{2p}.$ (72)
@@ I.5 Proof of Lemma 15
$F(\tilde{X}_{(t+1)K})-F(\tilde{X}_{tK})\leq\langle\nabla F(\tilde{X}_{tK}),\tilde{X}_{(t+1)K}-\tilde{X}_{tK}\rangle+\frac{L}{2}\|\tilde{X}_{(t+1)K}-\tilde{X}_{tK}\|^{2}\,.$ (75)
@@
Here $\bar{Y}_{t}=\frac{1}{\sqrt{K}}\sum_{i=0}^{K-1}Z_{tK+i}$ and $\Delta_{t}=-\sum_{i=0}^{K-1}H_{t,i}(\nabla F(\hat{\tilde{X}}_{tK+i})-\nabla F(\tilde{X}_{tK}))$.
Thus, summing Equation (75) from $t=0$ to t=T1t=T-1, we conclude:
@@
$\leq\sum_{t=0}^{T-1}-\tfrac{\alpha}{4}\|\nabla F(\tilde{X}_{tK})\|^{2}+3\alpha\|\Delta_{t}\|^{2}+3L\alpha\|\bar{Y}_{t}\|^{2}+\sqrt{2\alpha}\langle\nabla F(\tilde{X}_{tK}),\bar{Y}_{t}\rangle\,.$ (76)
@@
$\quad+\frac{C_{p}}{\alpha^{\frac{p}{2}}}\bigr|\sum_{t=0}^{T-1}\langle\nabla F(\tilde{X}_{tK}),\bar{Y}_{t}\rangle\bigr|^{p}\,.$ (77)
@@
$\leq\frac{\kappa C_{p}}{2\alpha^{p}}+\frac{C_{p}}{2\kappa}\mathbb{E}(\sum_{t=0}^{T-1}\|\nabla F(\tilde{X}_{tK})\|^{2})^{p}\,.$ (78)
@@
$\leq C_{p}T^{p-1}\sum_{t=0}^{T-1}L^{2p}\mathbb{E}(\sup_{0\leq i\leq K-1}\|\hat{\tilde{X}}_{tK+i}-\tilde{X}_{tK}\|)^{2p}\,.$ (79)
In the first step, we have used Jensen’s inequality. In the second step we have used the fact that $\nabla F$ is $L$-Lipshcitz and that $N_{t}$ is independent of $\hat{\tilde{X}}_{tK+i}$ for $0\leq i\leq K-1$. In the last step, we have used the fact that $\mathbb{E}(N_{t})^{2p}\leq C_{p}$ for some constant $C_{p}$ (Lemma 28).
@@
$\mathbb{E}\sup_{0\leq i\leq k}\|\hat{\tilde{X}}_{tK+i}-\tilde{X}_{tK}\|^{2p}\leq C_{p}\alpha^{2p}\mathbb{E}\|\nabla F(\tilde{X}_{tK})\|^{2p}+C_{p}\alpha^{p}d^{p}\,.$
@@
$\quad+C_{p}L^{p}d^{p}T^{p}+\frac{C_{p}}{\alpha^{p}}\,.$ (80)
@@ I.8 Proof of Lemma 21
    Using the scaling relations (Section 2.1), it is easy to show that for any $i$ such that $0\leq i\leq K-1$, $\sum_{j=0}^{i-1}\Pi A_{\tfrac{\alpha(i-1-j)}{K}}\Gamma_{\tfrac{\alpha}{K}}Z_{tK+j}$ is a Gaussian with covariance matrix $\Sigma_{i}$ such that $\Sigma_{i}\preceq\frac{2\exp(\gamma\alpha)\gamma\alpha^{3}}{3}\mathbf{I}_{d}$.
    Applying Gaussian concentration for $\sum_{j=0}^{i-1}\Pi A_{\tfrac{\alpha(i-1-j)}{K}}\Gamma_{\tfrac{\alpha}{K}}Z_{tK+j}$ along with the union bound over all $0\leq i\leq K-1$, we conclude:
@@ I.9 Proof of Lemma 22
    In the second step, we have use the fact that $|\frac{2g(1-g)}{\gamma}\langle\nabla F(\tilde{U}_{tK}),\tilde{V}_{tK}\rangle|\leq g(1-g)\frac{\|\nabla F(\tilde{U}_{tK})\|^{2}}{\gamma^{2}}+g(1-g)\|\tilde{V}_{tK}\|^{2}$. In the third step, we have used the fact that: $2g\langle\tilde{V}_{tK},(\mathbf{I}-\Pi)\alpha\Delta_{t}\rangle\leq(\sqrt{g}-g)\|\tilde{V}_{tK}\|^{2}+\frac{\|(\mathbf{I}-\Pi)\alpha\Delta_{t}\|^{2}}{1-\sqrt{g}}$, $\frac{2(1-g)}{\gamma}\langle\nabla F(\tilde{U}_{tK}),(\mathbf{I}-\Pi)\alpha\Delta_{t}\rangle\leq\frac{1-g}{\gamma^{2}}\|\nabla F(\tilde{U}_{tK})\|^{2}+(1-g)\|(\mathbf{I}-\Pi)\alpha\Delta\|^{2}$ and $\|(\mathbf{I}-\Pi)(\alpha\Delta_{t}+\Gamma_{\alpha}\tilde{Y}_{t})\|^{2}\leq 2\|(\mathbf{I}-\Pi)(\alpha\Delta_{t})\|^{2}+2\|(\mathbf{I}-\Pi)\Gamma_{\alpha}\tilde{Y}_{t}\|^{2}$.
@@
    We now use the fact that for any two random variables $X,Y$ such that $\mathbb{E}\|X\|^{p},\mathbb{E}\|Y\|^{p}<\infty$, $\mathbb{E}\|X+Y\|^{p}]^{\frac{1}{p}}\leq[\mathbb{E}\|X\|^{p}]^{\frac{1}{p}}+[\mathbb{E}\|Y\|^{p}]^{\frac{1}{p}}$. Using the bounds established above and applying them to Equation (91), we conclude:
@@
    Summing item 2 of Lemma 21 over $t=0$ to t=T1t=T-1 and applying the triangle inequality for $p$-th moments, we have: